OpenStudy (anonymous):

**MATH HELP!!** WILL FAN AND GIVE A MEDAL!!!*** Please i need to have this done in 1 hour!

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

that screenshot :)

OpenStudy (anonymous):

Ah, a test. I'll let someone else do your work for you.

OpenStudy (anonymous):

5 is option B because you divide the total price by the number of bagels.

OpenStudy (anonymous):

If one gallon cots 1.94 then 8 must cost 1.94 * 8

OpenStudy (anonymous):

1,94* 8 = $15.52

OpenStudy (anonymous):

Question 6 is option A because you compare the price per unit with one other.

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

And question 7 is option D, you divide each fraction by your given fraction.

OpenStudy (anonymous):

4. 1 gallon cost $1.94, then 8 gallons cost $1.94*8= $15.52 5. 7 bagels cost $2.94, then one bagel costs $2.94/7=$0.42 6. A, because it has the lowest unit cost $10.88/7=$1.55 7. D because 8*9 = 72, and 3*9 = 27 therefore 72/27 = 8/3